# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v2

@@ 3 DiverseHead (Cross-head Supervision)
Figure 2: DiverseHead: an online semi-supervised learning approach. This figure applies the dynamic freezing strategy: slide windows of the freezer randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation). Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet only a single head is randomly chosen to be updated by unsupervised loss.
@@
This section introduces the details of the proposed DiverseHead semi-supervised learning method which is proposed to create various predictions by multiple heads illustrated in Figure 2, specifically the number of convolutional layers for each head is 2. The source of perturbation is from two strategies, called dynamic freezing and dropout which are shown in Figure 4 and explained in detail in Section 3.1. The proposed DiverseHead does not distinguish between the main or auxiliary decoder, which differs from CCT. During every training iteration, all heads undergo supervision through supervised losses, yet only a single head is randomly chosen to be updated by an unsupervised loss. In addition, the proposed method combines self-training and consistency regularization, which uses their predictions to supervise itself and forces all perturbed head to give a consistent output. It makes an efficient voting module shown in Figure 3 to help calculate unsupervised loss by using optimized pseudo labels. To specifically explain the proposed semi-supervised framework, given both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads. Each of these heads is initialized differently. The expectation of the proposed semi-supervised learning architecture is to gain the trained network $Q$ leveraging these labelled and unlabelled data.
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k^{\prime}th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{j},y_{k}\right),$ (1)
where $W$ and $H$ refer to the weight and height of input images. The final supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}$ (2)
@@
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head. Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels. There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively. The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$. Subsequently, this combined prediction is used to calculate the mean pseudo labels $\hat{y}_{k}^{mean}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Apart from the mean voting module, the individual pseudo label is generated from the output of each head by the $argmax$ function. The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight. $\varphi$ is a learnable parameter and its value changes depending on the dataset and training processing. The max voting module returns the class number that gets the most votes for each pixel. After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss by using the cross-entropy loss function:
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{final}\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head. There is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss.
Finally, the whole loss is written as
$\mathcal{L}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup},$ (4)
where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
(a) Dynamic Freezing
(b) Dropout
Figure 4: The proposed two perturbation methods: dropout and dynamic freezing. In (a), the dynamic selector is designed to amplify the parameter diversity of multiple heads. It randomly chooses a set number of heads not to be updated by backpropagation during each training iteration, thereby enhancing the variability of parameters.
3.1 Perturbation Methods
Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
@@ 3.1 Perturbation Methods
              Initialise model $Q$
              Randomly initialise parameters for each head
@@
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
@@
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
@@
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
        $\mbox{\it Unfreeze}(\left\{head^{i}\right\}_{i=R})$
@@
Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
@@
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
@@
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
@@
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
@@
Another form of perturbation involves the use of dropout layers to the proposed DiverseHead structure to increase the diversity of features during training. We introduce a dropout layer after each convolutional block in each segmentation head of the network as shown in Figure 4 (b). The pseudocode of this method is given in Algorithm 2. As both dynamic freezing and dropout perturbation utilize the same semi-supervised learning framework, DiverseHead, the network initialization and supervised loss calculation procedures stay the same. Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section 6. Considering various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
4 Cross-model Supervision (DiverseModel)
@@
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks. In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet (Zhao et al., 2017), UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017). Algorithm 3 presents the pseudocode of the DiverseModel method. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020) technique for every network employed within the framework of the DiverseModel architecture. Grad-CAM visualizes the areas of an image that are important to the model predictions from each network. Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set. Examining Figure 6, we can see that different networks pay different and complementary attention to the same input, and the pseudo labels from the DiverseModel prediction show the highest quality.
@@
Figure 6: The upper section displays Grad-CAM outputs from individual networks within the DiverseModel architecture using the Potsdam dataset. The lower section showcases both the ground truth and pseudo labels generated through predictions from each network. In the lower right corner, the pseudo label from DiverseModel is presented.
@@
The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
$\mathcal{L}_{sup}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{n},y_{k}\right),$ (5)
@@
where $p_{k}^{n}$ represents the prediction of the $k\leavevmode\nobreak\ th$ pixel from the $n\leavevmode\nobreak\ th$ network.
@@
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in (Ma et al., 2023), all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right).$ (6)
Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
@@
               Randomly initialise 3 models, PSPNet $Q^{1}$, UNet $Q^{2}$, SegNet $Q^{3}$
@@
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
                  + $loss(Q^{3}(x_{k}),y_{k})$ based on (1)
@@
        $\mathcal{L}_{unsup}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
@@
        $\mathcal{L}\leftarrow\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
@@
The cross-pseudo supervision among three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\frac{1}{6}\left(\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}+\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}\right).\end{array}$ (7)
@@
The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
5 Dataset Description

### v3

@@ III DiverseHead (Cross-head Supervision)
Figure 2: DiverseHead: an online semi-supervised learning approach. This figure applies the dynamic freezing strategy: the freezers (Dynamic Selector in the figure) randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation). Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet each head is randomly chosen to be updated by an unsupervised loss.
@@
This section introduces the details of the proposed DiverseHead SSL method, which uses a single network with multiple heads, illustrated in Figure 2. Each head has 2 convolutional layers. Unlike CCT, the proposed DiverseHead method treats all heads equally, avoiding a distinction between the main and auxiliary heads. Meanwhile, all heads benefit from labelled data by applying supervised losses to each. This helps create better pseudo labels in the subsequent steps of the proposed method. On the other hand, rather than adding artificial noise to perturb head features as in CCT, DiverseHead introduces two perturbation strategies: Dynamic Freezing and Dropout. They are demonstrated in Figure 4 and explained in detail in Section III-A. During each training iteration, in addition to supervised losses, an unsupervised loss is computed between the pseudo-label and the prediction. The pseudo label is derived from an efficient voting module illustrated in Figure 3. Although non-differentiable operations (e.g. Voting and Argmax) could potentially limit optimisation during training, in our framework, parameter perturbations, feature perturbations, and independent initialisation across heads and models inject sufficient diversity and stochasticity into the training process. These factors enable effective gradient-based optimisation despite the presence of non-differentiable components. The proposed method can be seen as a combination of self-training and consistency regularisation, leveraging model predictions to supervise itself and forcing all perturbed heads to produce consistent outputs.
To provide a detailed description of the proposed semi-supervised framework, given both a labelled data set $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads. Each of these heads is initialised differently. The proposed SSL approach expects to gain the trained network $Q$ leveraging the labelled and unlabelled data.
The output of each head refers to a probability-based prediction for all classifications. When working with labelled data, the supervised loss $\mathcal{L}_{sup,s}^{j}$ between the $s^{\prime}th$ ground truth $y_{s}$ and its corresponding prediction $p_{s}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup,s}^{j}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,s}^{j},y_{i,s}\right),$ (1)
where $W$ and $H$ refer to the width and height of input images. The final supervised loss for the $s^{\prime}th$ labelled data is determined by the mean of the losses across all heads
$\mathcal{L}_{sup,s}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}$ (2)
@@
In each iteration, we also get a prediction $r_{k}^{j}$ corresponding to the $k^{\prime}th$ unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head. Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels. There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively shown in Figure 3. The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$. Subsequently, this combined prediction is used to calculate the mean pseudo labels $\hat{y}_{k}^{mean}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Apart from the mean voting module, the individual pseudo label is generated from the output of each head by the $argmax$ function. The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight. $\varphi$ is a learnable parameter and its value changes depending on the dataset and training process. The max voting module returns the class number that gets the most votes for each pixel. After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss $\mathcal{L}_{unsup,k}$ by using the cross-entropy loss function:
$\mathcal{L}_{unsup,k}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(r_{i,k}^{main},\hat{y}_{i,k}^{final}\right),$ (3)
where $r_{k}^{main}$ is the prediction from the randomly selected single head.
Finally, the whole loss can be written as
$\mathcal{L}=\mathcal{L}_{sup,s}+\lambda\mathcal{L}_{unsup,k},$ (4)
where $\lambda$ is the trade-off weight between the supervised and unsupervised losses. We use different subscripts ($s$ and $k$) for the supervised and unsupervised losses, respectively, to account for the difference in the numbers of labelled and unlabelled data. Specifically, the number of labelled data samples is smaller than that of unlabelled data samples. The solution to this issue is that, in each epoch, the labelled data set is repeatedly used in cycles for multiple iterations until all unlabelled data have been processed once.
Figure 4: The proposed two perturbation methods: (A) Dynamic Freezing and (B) Dropout. Dynamic Freezing was designed to enhance the parameter diversity across multiple heads. During each training iteration, a specific subset of heads is randomly selected (The Dynamic Selector in the figure is used for this purpose), and their parameters are frozen, meaning they are not updated by minimising the loss in that iteration. These parameters are unfrozen before the next iteration begins. Instead, for Dropout, each channel of features passed through each head is independently zeroed out with a dropout rate $p$ during each forward pass.
III-A Perturbation Methods
Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode. The Labelled Training Dataset is defined as $\mathcal{B}^{l}$. Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch. We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
@@ III-A Perturbation Methods
  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
@@
             $p_{s}^{j}=Q^{head_{j}}(x_{s}^{j})$
             $\mathcal{L}_{sup,s}^{j}=loss(p_{s}^{j},y_{s})$ based on (1)
@@
        $\mathcal{L}_{unsup,k}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}+\lambda\mathcal{L}_{unsup,k}$
        Minimize $\mathcal{L}$ to update model $Q$
        $\mbox{{Unfreeze}}(\left\{head^{i}\right\}_{i=R})$
@@
Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode. The Labelled Training Dataset is defined as $\mathcal{B}^{l}$. Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch. We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
@@
  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
@@
             $p_{s}^{j}=Q^{head_{j}}(x_{s}^{j})$
             $\mathcal{L}_{sup,s}^{j}=loss(p_{s}^{j},y_{s})$ based on (1)
@@
        $\mathcal{L}_{unsup,k}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}+\lambda\mathcal{L}_{unsup,k}$
        Minimize $\mathcal{L}$ to update model Q
@@
Another form of perturbation involves the use of dropout layers in the proposed DiverseHead structure to increase the diversity of features during training. We introduce a dropout layer after each convolutional block in each segmentation head of the network, as shown in Figure 4 (b). The pseudocode of this method is given in Algorithm 2. As both dynamic freezing and dropout perturbation utilise the same semi-supervised learning framework, DiverseHead, the network initialisation and supervised loss calculation procedures stay the same. Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.5. Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section VI. Considering that various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
IV Cross-model Supervision (DiverseModel)
@@
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks. In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet [35], UNet [34], SegNet [33]. Algorithm 3 presents the pseudocode of the DiverseModel method. Since different networks pay different and complementary attention to the same input, this offers the basis for they are able to benefit from each other. In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) [26] technique for every network employed within the framework of the DiverseModel architecture. Grad-CAM visualises the areas of an image that are important to the model predictions from each network. Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set. Examining Figure 6, we can see that different networks pay different and complementary attention to the same input, and the pseudo labels from the DiverseModel prediction show the highest quality.
@@
Figure 6: The upper section displays Grad-CAM outputs from individual networks within the DiverseModel architecture using the Potsdam dataset. The lower section showcases both the ground truth and pseudo labels generated through predictions from each network. In the lower right corner, the pseudo-label from DiverseModel is presented.
@@
The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss $\mathcal{L}_{sup,s}$ is expressed as:
$\mathcal{L}_{sup,s}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,s}^{n},y_{i,s}\right),$ (5)
@@
where $p_{i,s}^{n}$ represents the $i\leavevmode\nobreak\ th$ predicted pixel of the $s\leavevmode\nobreak\ th$ sample from the $n\leavevmode\nobreak\ th$ network.
@@
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in [25], all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup,k}^{12}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,k}^{1},\hat{r}_{i,k}^{2}\right).$ (6)
where, $p_{i,k}^{1}$ represents the $i\leavevmode\nobreak\ th$ predicted pixel of the $k\leavevmode\nobreak\ th$ sample from the $1\leavevmode\nobreak\ st$ network.
Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode. The Labelled Training Dataset is defined as $\mathcal{B}^{l}$. Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch. We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
@@
                Randomly initialise three models, PSPNet $Q^{1}$,
                UNet $Q^{2}$, SegNet $Q^{3}$
@@
  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        $\mathcal{L}_{sup,s}=loss(Q^{1}(x_{s}),y_{s})$ + $loss(Q^{2}(x_{s}),y_{s})$
                    + $loss(Q^{3}(x_{s}),y_{s})$ based on (1)
@@
        $\mathcal{L}_{unsup,k}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
@@
        $\mathcal{L}\leftarrow\mathcal{L}_{sup,s}+\lambda\mathcal{L}_{unsup,k}$
        Minimize $\mathcal{L}$ to update model $Q^{1}$, $Q^{2}$, $Q^{3}$
@@
The cross-pseudo supervision among the three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
$\mathcal{L}_{unsup,k}$ $=\frac{1}{6}\Big{(}\mathcal{L}_{unsup,k}^{12}+\mathcal{L}_{unsup,k}^{13}+\mathcal{L}_{unsup,k}^{21}$ (7)
$\quad+\mathcal{L}_{unsup,k}^{23}+\mathcal{L}_{unsup,k}^{31}+\mathcal{L}_{unsup,k}^{32}\Big{)}.$
@@
The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup,s}$ and $\mathcal{L}_{unsup,k}$, which is previously given in equation (4)
V Dataset Description